Assuming earth to be spherical, at what height above the north pole, value of g is same as that on the surface of earth at equator.

Grade:11

5 Answers

Anushka Prasad
11 years ago

Since it is a spherical radius will be same everywhere.

Therefore at zero height above the north pole value of g is same as that on the surface of earth at equator. (h=0)
 

 Smile

Arnab Mandal
11 years ago

Answer is given 10 km.

nikhil manali
11 years ago

intelligent ans man

Nishant Singh
5 years ago
g at height h = g at equator. (acc. to question)g(1-2h/R) = g - ß²R. (ẞ here represents OMEGA)g- 2gh/R = g - ß²ROn solving,2gh/R = ß²R2gh = 0.034 × R ( Omega²R = 0.034 , STANDARD VALUE)ON SOLVING, h= 34×32/100h = 10.88 Km ✓ is the correct Answer
Nishant Singh
5 years ago
g at height h = g at equator. (acc. to question)g(1-2h/R) = g - ß²R. (ẞ here represents OMEGA)g- 2gh/R = g - ß²ROn solving,2gh/R = ß²R2gh = 0.034 × R ( Omega²R = 0.034 , STANDARD VALUE)ON SOLVING, h= 34×32/100h = 10.88 Km ✓